OpenStudy (anonymous):

Help with Trigonomic ratios!!! I have a right triangle with another angle at 51 degrees and the opposite side is 6 how do I find the length of the other leg?

OpenStudy (dumbcow):

SOH CAH TOA the 2 legs are the opposite and adjacent sides, so we need to use the tangent ratio TOA ---> tan(51) = opp/adj = 6/x solve for x

OpenStudy (anonymous):

would it be 7.409?

OpenStudy (anonymous):

I've tried to do these kinds of equations on a few tests so far and always get them wrong

OpenStudy (dumbcow):

no \[\frac{6}{x} = \tan(51)\] flip both sides \[\frac{x}{6} = \frac{1}{\tan(51)}\] multiply by 6 \[x = \frac{6}{\tan(51)}\]

OpenStudy (anonymous):

ohhhhh okay. so it would be 4.858?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Okay. Could you help me with one more? It's probably easy like this one but I'm just not getting these...

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

The problem is tan(x)=5/1 and my course never really taught me this…

OpenStudy (dumbcow):

ok in this case we want to find the angle because the variable is inside the tangent function, the only way to solve for "x" is to get rid of tangent by using the inverse tangent same reason you get rid of x^2 by using square root \[\tan^{-1} (\tan x) = \tan^{-1} (\frac{5}{1})\] \[x = \tan^{-1} (5)\] use your calculator, make sure you are in degree mode

OpenStudy (anonymous):

78.690?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Yay! Thank you so much!!

OpenStudy (dumbcow):

yw :) by the way, are you from phoenix?

OpenStudy (anonymous):

No haha. I just really love mythology and Phoenixlizzie was my username from another site. My favorite mythological creature and my middle name!

OpenStudy (dumbcow):

ok cool .... i live in phoenix

OpenStudy (anonymous):

Oh cool! I'm from NH but we just moved to SC. The heat is killing me here- I couldn't imagine Phoenix!